OpenStudy (anonymous):

HOw to solve in paper plz : \[\sqrt{7} \sqrt{28}\]

OpenStudy (anonymous):

the answer is 14. Multiply the \[\sqrt{7} \sqrt{28}\] then you get 14

OpenStudy (anonymous):

\[\sqrt{7}\sqrt{28} = \sqrt{7} \sqrt{4 \times 7} = \sqrt{7}\sqrt{7}\sqrt{4} = 7 \sqrt{4} = 14\]

OpenStudy (anonymous):

OKie Thank you :D I knew it was 14 be coz of the calcu but wated to learn without the calcu. TY so much . Do u play redstone